# Wonder3D++: Cross-domain Diffusion for High-fidelity 3D Generation from a Single Image

### v1

@@ I Introduction
Reconstructing 3D geometry from a single image [25, 46, 13, 36, 42, 49, 43] stands as a fundamental task in computer graphics and 3D computer vision, benefiting a wide range of versatile applications such as novel view synthesis [39, 24, 7], 3D content creation [38, 53], and robotics grasping [93, 29]. However, this task is notably challenging since it is ill-posed and demands the ability to discern the 3D geometry of both visible and invisible parts. This ability requires extensive knowledge of the 3D world.
@@
Recently, the field of 3D generation has experienced rapid and flourishing development with the introduction of diffusion models. A growing body of research [52, 71, 74, 34, 5], such as DreamField [23], DreamFusion [52], and Magic3D [34], resort to distilling prior knowledge of 2D image diffusion models or vision language models to create 3D models from text or images via Score Distillation Sampling (SDS) [52]. Despite achieving some degree of generalization performance, these methods still encounter challenges like multi-view inconsistencies and low-quality results.
@@
To enhance the multi-view consistency, several methods emerged that directly generate multi-view 2D images, with representative works including SyncDreamer [38] and MVDream [60]. By improving information exchange between views using attention mechanisms, existing methods generate multi-view images and then optimize 3D representations—such as Neus [72] and NeRF [47]—to reconstruct textured meshes. However, current multi-view generative methods face significant challenges. First, relying solely on color images often compromises the fidelity of the generated shapes, making it difficult to recover geometric details without incurring high computational costs. Additionally, most multi-view diffusion models are constrained to predefined camera types, which can introduce distortions when processing images captured in uncontrolled environments, further amplifying multi-view inconsistencies.
@@
Recently, various techniques have been introduced that aim to generate 3D geometries via network inference, including point clouds [49, 88, 45, 94], meshes [40, 16], neural fields [73, 8, 26, 28, 17, 1, 48, 50, 25, 89, 19, 15, 4, 21]. Most of them attempt to train 3D generative models from scratch on datasets consisting of various 3D assets. To further enhance generalizability, recent approaches have focused on generating sparse view images using a multi-view diffusion model as a supplementary cue, followed by a feed-forward sparse-view reconstruction model to create 3D objects [63, 75, 83, 82]. However, despite some of these methods requiring over a hundred GPUs for training, they still face significant challenges in achieving high-quality geometry and textures. Additionally, since these methods rely on neural networks for direct forward propagation to generate 3D models, the resulting objects often lack precise alignment with the input images.
@@
To address the challenges of fidelity, consistency, and generalizability in previous works, we propose a novel approach for single-view 3D reconstruction by generating multi-view consistent normal maps and their corresponding color images using a cross-domain diffusion model. The core idea is to extend diffusion frameworks to jointly model the distribution of two domains: normals and colors. Normal maps capture surface undulations and variations, encoding detailed geometric information that allows for high-fidelity 3D geometry extraction from 2D representations. By leveraging 2D representations, Wonder3D++ is built on the pre-trained Stable Diffusion model [54], utilizing strong priors that enable robust zero-shot generalization.
@@
We conduct extensive experiments on the Google Scanned Object dataset [14] and various 2D images with different styles. The experiments validate that Wonder3D++ achieves a leading level of geometric details with high efficiency and strong generalization among current zero-shot single-view reconstruction methods.
Figure 2: Overview of Wonder3D++. Given a single image, Wonder3D++ takes the input image feature produced by VAE encoder, the image embedding produced by CLIP model [54], the camera parameters of multiple views, a camera type switcher and a domain switcher as conditioning to generate consistent multi-view normal maps and color images. Subsequently, Wonder3D++ applies an innovative cascaded 3D mesh reconstruction algorithm, which utilizes a coarse-to-fine strategy consisting of geometric initialization, coarse mesh reconstruction, and iterative mesh refinement to produce a 3D mesh with high-quality geometry and high-fidelity textures.
Differences with conference version. Wonder3D++ introduces several key improvements over Wonder3D [41], significantly enhancing robustness and quality: 1) the addition of a camera type switcher, allowing the model to handle input images captured with various camera settings; 2) the implementation of multi-domain pre-training, ensuring a seamless transition from stable diffusion pre-training to the desired joint distribution for 3D data; and 3) replacing implicit neural SDF fields with explicit mesh representations for 3D textured mesh reconstruction, introducing a cascaded 3D mesh extraction algorithm that greatly improves both efficiency and effectiveness.
@@ II Related Works
Recently, benefiting from the availability of large-scale high-quality data, there have been significant advancements in foundational models, such as 2D diffusion models [20, 56, 9] and large vision-language models (e.g., the CLIP model [54]). These models offer new opportunities for generating 3D assets by leveraging the strong priors provided by these pre-trained architectures. In comparison, directly trained native 3D generators often struggle to ensure both generalizability and quality, due to the relative scarcity of large-scale data in the three-dimensional domain. Therefore, pioneering works such as DreamFusion [52] and SJC [71] have proposed distilling a 2D text-to-image generation model to generate 3D shapes from textual descriptions, and subsequent works have adopted this per-shape optimization scheme, building upon the success of these initial approaches. For the task of text-to-3D [5, 74, 57, 86, 34, 58, 67, 96, 22, 2, 78, 6] or image-to-3D synthesis [64, 46, 53, 81, 55, 59], these methods typically optimize a 3D representation (i.e., NeRF, mesh, or SDF), and then leverage differentiable rendering to deliver 2D images from various viewpoints. The images are then fed into the 2D diffusion models or CLIP model for calculating SDS [52] losses, which guide the 3D shape optimization. While this endows 3D generation methods with unprecedented generalization capabilities, most of these methods suffer from low efficiency and the “multi-face” problem (i.e., the Janus problem [52]) due to the lack of explicit 3D supervision. Moreover, per-shape optimization, which typically takes tens of minutes, significantly limits the practicality of this methodology.
@@ II-B Multi-view Diffusion Models
To alleviate the ”multi-face” problem, some efforts [77, 18, 11, 95, 68, 3, 87, 66, 85, 62, 65, 80, 37, 31, 90] are made to extend 2D diffusion models from single-view images to multi-view consistent images. However, most of these methods focus on image generation and are not designed for 3D reconstruction. The recent works Viewset Diffusion [62], SyncDreamer [38], and MVDream [60] share a similar idea to produce consistent multi-view color images via attention layers. However, unlike that normal maps explicitly encode geometric information, reconstruction from color images always suffers from texture ambiguity, and, thus, they either struggle to recover geometric details or require huge computational costs. For instance, SyncDreamer [38] requires dense views for 3D reconstruction but still suffers from low-quality geometry and blurring textures. MVDream [60] still resorts to a time-consuming optimization using SDS loss for 3D reconstruction, and its multi-view distillation scheme requires 1.5 hours.
Direct2.5[44] employs a sequential process: it first generates multi-view normal maps and subsequently generates RGB images. Consequently, the generation of normal maps and RGB images is largely decoupled. This separation can be less conducive to achieving strong alignment between the geometric and textural domains, or to effectively learning a coherent 3D data distribution. Moreover, the inconsistency between the camera type of the input images and the predefined camera type leads to unnatural distortions in the generated meshes. In contrast, thanks to our proposed domain switcher, our method can simultaneously generate multi-view normal maps for geometric reconstruction and RGB images for coloring. Additionally, our method provides the flexibility to generate results in either orthogonal or perspective camera settings, adapting to different requirements as needed.
@@
Instead of performing a time-consuming per-shape optimization guided by 2D diffusion models, some works attempt to directly train 3D diffusion models based on various 3D representations, like point clouds [49, 88, 45, 94], meshes [40, 16], neural fields [73, 8, 26, 28, 17, 1, 48, 50, 25, 89, 19, 15, 4]. However, due to the limited size of publicly available 3D assets datasets, most of the works have only been validated on limited categories of shapes or lack of sufficient generalization properties, and how to scale up on large datasets is still an open problem. In order to effectively improve the generalization performance of the model, a portion of recent work [75, 82, 63, 83, 33] seeks to first utilize the multi-view diffusion model to infer multi-view images and then use sparse view reconstruction models to obtain meshes in a single forward pass. Such sparse view reconstruction models are usually based on transformer architectures to predict Nerf [82], Gaussian Splatting [63, 83] or triplane [75] representations and employ differentiable rendering to compute the loss function, which hinders the efficiency and stability of the training process. As a result, the training process typically requires tens or hundreds of GPUs and it is difficult to obtain geometries and textures with sufficient details. On the contrary, our method first adopts 2D representations, enabling it to leverage 2D diffusion models [56], whose pre-trained priors significantly enhance zero-shot generalization ability. Building on this, we then employ a cascaded approach to efficiently reconstruct 3D meshes, improving both geometric and texture quality while avoiding the high computational costs associated with training a feedforward reconstruction model.
@@ III-A Diffusion Models
Diffusion models [20, 61] are first proposed to gradually recover images from a specifically designed degradation process, where a forward Markov chain and a Reverse Markov chain are adopted. Given a sample $z_{0}$ drawn from the data distribution $p(z)$, the forward process of denoising diffusion models yields a sequence of noised data $\left\{z_{t}\mid t\in(0,T)\right\}$ with $z_{t}=M_{t}z_{0}+\sigma_{t}\epsilon$, where $\epsilon$ is random noise drawn from distribution $\mathcal{N}(0,1)$, and $M_{t},\sigma_{t}$ are fixed sequence of the noise schedule. The forward process will be iteratively applied to the target image $z$ until the image becomes complete Gaussian noise at the end. On the contrary, the reverse chain iteratively denoises the corrupted image, recovering $z_{t-1}$ from $z_{t}$ by step-by-step predicting the added random noise $\epsilon$.
@@ III-C Mesh Reconstruction via Differentiable Rendering
Given the synthesized multi-view normal maps $N^{1:K}$, we extract the 3D mesh using differentiable rendering, ensuring accurate geometry reconstruction through gradient-based optimization. Specifically, starting with an initial triangle mesh $M$, composed of vertices $V$ and faces $F$, the multi-view normal maps $N^{i:K}$ and masks $M^{i:K}$ are rendered differentiably using the nvdiffrast library [30]. The reconstruction loss is then computed, and the gradient is back-propagated to optimize the vertices $V$, refining the mesh geometry based on the multi-view input.
To directly optimize the vertices of the mesh and continuously adjust the topology of the mesh and the edge lengths of the mesh, we adopt the remeshing strategy proposed by  [51], and apply it after each optimization iteration. This process involves edge-splitting, edge-collapsing, edge-flipping, and smoothing to adaptively adjust local edge lengths and correct defects caused by sparse or inconsistent vertex gradients.
The reasons we choose an explicit mesh representation over implicit methods like Neus [72] are twofold: 1) The computational complexity of optimizing with the explicit method scales with the number of mesh vertices and the square of the resolution of the multi-view normal maps, whereas in implicit neural field methods, it scales with the cube of the image resolution, making our approach more efficient. 2) Directly optimizing mesh vertices and faces allows our method to integrate seamlessly with the cascaded mesh extraction process, eliminating the computational overhead and detail loss that typically occur during conversions between mesh and implicit SDF representations.
@@ IV-A Multi-view Cross-domain Diffusion Model
Our model is built upon the pre-trained stable diffusion models [54] to leverage its strong priors. Given that current 2D diffusion models [54, 36] are designed for a single domain, the main challenge lies in how to effectively extend stable diffusion models to support multi-domain operations. To achieve this goal, we explore several possible designs.
@@ IV-A2 Domain Switcher
The domain switcher is first encoded via positional encoding [47] and subsequently concatenated with the time embedding. This combined representation is then injected into the UNet of the stable diffusion models. Interestingly, experiments show that this subtle modification does not significantly alter the pre-trained priors. As a result, it allows for fast convergence and robust generalization, without requiring substantial changes to the stable diffusion models.

### v2

@@ I Introduction
Reconstructing 3D geometry from a single image [jun2023shap, melas2023realfusion, dou2023tore, liu2023zero, Long_2021_ICCV, nichol2022point, long2024adaptive] stands as a fundamental task in computer graphics and 3D computer vision, benefiting a wide range of versatile applications such as novel view synthesis [liu2022neural, jiang2023gaussianshader, cheng2024gaussianpro], 3D content creation [liu2023syncdreamer, qian2023magic123], and robotics grasping [zheng2024gaussiangrasper, kleeberger2020survey]. However, this task is notably challenging since it is ill-posed and demands the ability to discern the 3D geometry of both visible and invisible parts. This ability requires extensive knowledge of the 3D world.
@@
Recently, the field of 3D generation has experienced rapid and flourishing development with the introduction of diffusion models. A growing body of research [poole2022dreamfusion, wang2023score, wang2023prolificdreamer, lin2023magic3d, chen2023fantasia3d], such as DreamField [jain2022zero], DreamFusion [poole2022dreamfusion], and Magic3D [lin2023magic3d], resort to distilling prior knowledge of 2D image diffusion models or vision language models to create 3D models from text or images via Score Distillation Sampling (SDS) [poole2022dreamfusion]. Despite achieving some degree of generalization performance, these methods still encounter challenges like multi-view inconsistencies and low-quality results.
@@
To enhance the multi-view consistency, several methods emerged that directly generate multi-view 2D images, with representative works including SyncDreamer [liu2023syncdreamer] and MVDream [shi2023mvdream]. By improving information exchange between views using attention mechanisms, existing methods generate multi-view images and then optimize 3D representations—such as Neus [wang2021neus] and NeRF [mildenhall2020nerf]—to reconstruct textured meshes. However, current multi-view generative methods face significant challenges. First, relying solely on color images often compromises the fidelity of the generated shapes, making it difficult to recover geometric details without incurring high computational costs. Additionally, most multi-view diffusion models are constrained to predefined camera types, which can introduce distortions when processing images captured in uncontrolled environments, further amplifying multi-view inconsistencies.
@@
Recently, various techniques have been introduced that aim to generate 3D geometries via network inference, including point clouds [nichol2022point, zeng2022lion, luo2021diffusion, zhou20213d], meshes [liu2023meshdiffusion, gao2022get3d], neural fields [wang2023rodin, cheng2023sdfusion, karnewar2023holofusion, kim2023neuralfield, gu2023learning, anciukevivcius2023renderdiffusion, muller2023diffrf, ntavelis2023autodecoding, jun2023shap, zhang20233dshape2vecset, gupta20233dgen, erkocc2023hyperdiffusion, chen2023single, hong2023lrm]. Most of them attempt to train 3D generative models from scratch on datasets consisting of various 3D assets. To further enhance generalizability, recent approaches have focused on generating sparse view images using a multi-view diffusion model as a supplementary cue, followed by a feed-forward sparse-view reconstruction model to create 3D objects [tang2024lgm, wang2024crm, xu2024grm, xu2024instantmesh]. However, despite some of these methods requiring over a hundred GPUs for training, they still face significant challenges in achieving high-quality geometry and textures. Additionally, since these methods rely on neural networks for direct forward propagation to generate 3D models, the resulting objects often lack precise alignment with the input images.
@@
To address the challenges of fidelity, consistency, and generalizability in previous works, we propose a novel approach for single-view 3D reconstruction by generating multi-view consistent normal maps and their corresponding color images using a cross-domain diffusion model. The core idea is to extend diffusion frameworks to jointly model the distribution of two domains: normals and colors. Normal maps capture surface undulations and variations, encoding detailed geometric information that allows for high-fidelity 3D geometry extraction from 2D representations. By leveraging 2D representations, Wonder3D++ is built on the pre-trained Stable Diffusion model [radford2021learning], utilizing strong priors that enable robust zero-shot generalization.
@@
We conduct extensive experiments on the Google Scanned Object dataset [downs2022google] and various 2D images with different styles. The experiments validate that Wonder3D++ achieves a leading level of geometric details with high efficiency and strong generalization among current zero-shot single-view reconstruction methods.
Figure 2: Overview of Wonder3D++. Given a single image, Wonder3D++ takes the input image feature produced by VAE encoder, the image embedding produced by CLIP model [radford2021learning], the camera parameters of multiple views, a camera type switcher and a domain switcher as conditioning to generate consistent multi-view normal maps and color images. Subsequently, Wonder3D++ applies an innovative cascaded 3D mesh reconstruction algorithm, which utilizes a coarse-to-fine strategy consisting of geometric initialization, coarse mesh reconstruction, and iterative mesh refinement to produce a 3D mesh with high-quality geometry and high-fidelity textures.
Differences with conference version. Wonder3D++ introduces several key improvements over Wonder3D [long2024wonder3d], significantly enhancing robustness and quality: 1) the addition of a camera type switcher, allowing the model to handle input images captured with various camera settings; 2) the implementation of multi-domain pre-training, ensuring a seamless transition from stable diffusion pre-training to the desired joint distribution for 3D data; and 3) replacing implicit neural SDF fields with explicit mesh representations for 3D textured mesh reconstruction, introducing a cascaded 3D mesh extraction algorithm that greatly improves both efficiency and effectiveness.
@@ II Related Works
Recently, benefiting from the availability of large-scale high-quality data, there have been significant advancements in foundational models, such as 2D diffusion models [ho2020denoising, rombach2022high, croitoru2023diffusion] and large vision-language models (e.g., the CLIP model [radford2021learning]). These models offer new opportunities for generating 3D assets by leveraging the strong priors provided by these pre-trained architectures. In comparison, directly trained native 3D generators often struggle to ensure both generalizability and quality, due to the relative scarcity of large-scale data in the three-dimensional domain. Therefore, pioneering works such as DreamFusion [poole2022dreamfusion] and SJC [wang2023score] have proposed distilling a 2D text-to-image generation model to generate 3D shapes from textual descriptions, and subsequent works have adopted this per-shape optimization scheme, building upon the success of these initial approaches. For the task of text-to-3D [chen2023fantasia3d, wang2023prolificdreamer, seo2023ditto, yu2023points, lin2023magic3d, seo2023let, tsalicoglou2023textmesh, zhu2023hifa, huang2023dreamtime, armandpour2023re, wu2023hd, chen2023it3d] or image-to-3D synthesis [tang2023make, melas2023realfusion, qian2023magic123, xu2022neurallift, raj2023dreambooth3d, shen2023anything], these methods typically optimize a 3D representation (i.e., NeRF, mesh, or SDF), and then leverage differentiable rendering to deliver 2D images from various viewpoints. The images are then fed into the 2D diffusion models or CLIP model for calculating SDS [poole2022dreamfusion] losses, which guide the 3D shape optimization. While this endows 3D generation methods with unprecedented generalization capabilities, most of these methods suffer from low efficiency and the “multi-face” problem (i.e., the Janus problem [poole2022dreamfusion]) due to the lack of explicit 3D supervision. Moreover, per-shape optimization, which typically takes tens of minutes, significantly limits the practicality of this methodology.
@@ II-B Multi-view Diffusion Models
To alleviate the ”multi-face” problem, some efforts [watson2022novel, gu2023nerfdiff, deng2023nerdi, zhou2023sparsefusion, tseng2023consistent, chan2023generative, yu2023long, tewari2023diffusion, yoo2023dreamsparse, szymanowicz2023viewset, tang2023mvdiffusion, xiang20233d, liu2023deceptive, lei2022generative, zhang2023text2nerf] are made to extend 2D diffusion models from single-view images to multi-view consistent images. However, most of these methods focus on image generation and are not designed for 3D reconstruction. The recent works Viewset Diffusion [szymanowicz2023viewset], SyncDreamer [liu2023syncdreamer], and MVDream [shi2023mvdream] share a similar idea to produce consistent multi-view color images via attention layers. However, unlike that normal maps explicitly encode geometric information, reconstruction from color images always suffers from texture ambiguity, and, thus, they either struggle to recover geometric details or require huge computational costs. For instance, SyncDreamer [liu2023syncdreamer] requires dense views for 3D reconstruction but still suffers from low-quality geometry and blurring textures. MVDream [shi2023mvdream] still resorts to a time-consuming optimization using SDS loss for 3D reconstruction, and its multi-view distillation scheme requires 1.5 hours.
Direct2.5[lu2024direct2] employs a sequential process: it first generates multi-view normal maps and subsequently generates RGB images. Consequently, the generation of normal maps and RGB images is largely decoupled. This separation can be less conducive to achieving strong alignment between the geometric and textural domains, or to effectively learning a coherent 3D data distribution. Moreover, the inconsistency between the camera type of the input images and the predefined camera type leads to unnatural distortions in the generated meshes. In contrast, thanks to our proposed domain switcher, our method can simultaneously generate multi-view normal maps for geometric reconstruction and RGB images for coloring. Additionally, our method provides the flexibility to generate results in either orthogonal or perspective camera settings, adapting to different requirements as needed.
@@
Instead of performing a time-consuming per-shape optimization guided by 2D diffusion models, some works attempt to directly train 3D diffusion models based on various 3D representations, like point clouds [nichol2022point, zeng2022lion, luo2021diffusion, zhou20213d], meshes [liu2023meshdiffusion, gao2022get3d], neural fields [wang2023rodin, cheng2023sdfusion, karnewar2023holofusion, kim2023neuralfield, gu2023learning, anciukevivcius2023renderdiffusion, muller2023diffrf, ntavelis2023autodecoding, jun2023shap, zhang20233dshape2vecset, gupta20233dgen, erkocc2023hyperdiffusion, chen2023single]. However, due to the limited size of publicly available 3D assets datasets, most of the works have only been validated on limited categories of shapes or lack of sufficient generalization properties, and how to scale up on large datasets is still an open problem. In order to effectively improve the generalization performance of the model, a portion of recent work [wang2024crm, xu2024instantmesh, tang2024lgm, xu2024grm, li2024craftsman] seeks to first utilize the multi-view diffusion model to infer multi-view images and then use sparse view reconstruction models to obtain meshes in a single forward pass. Such sparse view reconstruction models are usually based on transformer architectures to predict Nerf [xu2024instantmesh], Gaussian Splatting [tang2024lgm, xu2024grm] or triplane [wang2024crm] representations and employ differentiable rendering to compute the loss function, which hinders the efficiency and stability of the training process. As a result, the training process typically requires tens or hundreds of GPUs and it is difficult to obtain geometries and textures with sufficient details. On the contrary, our method first adopts 2D representations, enabling it to leverage 2D diffusion models [rombach2022high], whose pre-trained priors significantly enhance zero-shot generalization ability. Building on this, we then employ a cascaded approach to efficiently reconstruct 3D meshes, improving both geometric and texture quality while avoiding the high computational costs associated with training a feedforward reconstruction model.
@@ III-A Diffusion Models
Diffusion models [ho2020denoising, sohl2015deep] are first proposed to gradually recover images from a specifically designed degradation process, where a forward Markov chain and a Reverse Markov chain are adopted. Given a sample $z_{0}$ drawn from the data distribution $p(z)$, the forward process of denoising diffusion models yields a sequence of noised data $\left\{z_{t}\mid t\in(0,T)\right\}$ with $z_{t}=M_{t}z_{0}+\sigma_{t}\epsilon$, where $\epsilon$ is random noise drawn from distribution $\mathcal{N}(0,1)$, and $M_{t},\sigma_{t}$ are fixed sequence of the noise schedule. The forward process will be iteratively applied to the target image $z$ until the image becomes complete Gaussian noise at the end. On the contrary, the reverse chain iteratively denoises the corrupted image, recovering $z_{t-1}$ from $z_{t}$ by step-by-step predicting the added random noise $\epsilon$.
@@ III-C Mesh Reconstruction via Differentiable Rendering
Given the synthesized multi-view normal maps $N^{1:K}$, we extract the 3D mesh using differentiable rendering, ensuring accurate geometry reconstruction through gradient-based optimization. Specifically, starting with an initial triangle mesh $M$, composed of vertices $V$ and faces $F$, the multi-view normal maps $N^{i:K}$ and masks $M^{i:K}$ are rendered differentiably using the nvdiffrast library [laine2020modular]. The reconstruction loss is then computed, and the gradient is back-propagated to optimize the vertices $V$, refining the mesh geometry based on the multi-view input.
To directly optimize the vertices of the mesh and continuously adjust the topology of the mesh and the edge lengths of the mesh, we adopt the remeshing strategy proposed by  [palfinger2022continuous], and apply it after each optimization iteration. This process involves edge-splitting, edge-collapsing, edge-flipping, and smoothing to adaptively adjust local edge lengths and correct defects caused by sparse or inconsistent vertex gradients.
The reasons we choose an explicit mesh representation over implicit methods like Neus [wang2021neus] are twofold: 1) The computational complexity of optimizing with the explicit method scales with the number of mesh vertices and the square of the resolution of the multi-view normal maps, whereas in implicit neural field methods, it scales with the cube of the image resolution, making our approach more efficient. 2) Directly optimizing mesh vertices and faces allows our method to integrate seamlessly with the cascaded mesh extraction process, eliminating the computational overhead and detail loss that typically occur during conversions between mesh and implicit SDF representations.
@@ IV-A Multi-view Cross-domain Diffusion Model
Our model is built upon the pre-trained stable diffusion models [radford2021learning] to leverage its strong priors. Given that current 2D diffusion models [radford2021learning, liu2023zero] are designed for a single domain, the main challenge lies in how to effectively extend stable diffusion models to support multi-domain operations. To achieve this goal, we explore several possible designs.
@@ IV-A2 Domain Switcher
The domain switcher is first encoded via positional encoding [mildenhall2020nerf] and subsequently concatenated with the time embedding. This combined representation is then injected into the UNet of the stable diffusion models. Interestingly, experiments show that this subtle modification does not significantly alter the pre-trained priors. As a result, it allows for fast convergence and robust generalization, without requiring substantial changes to the stable diffusion models.